# Family Unification in a Six Dimensional Theory with an Orthogonal Gauge Group

### v1

@@
In this letter, we propose a new simple model of family unification avoiding the above unsatisfactory points. This model is also based on the GGHU, but an orthogonal group is employed as a gauge group. As was discussed in [25], the reason of considering the orthogonal group is that a spinor representation in the orthogonal group has a very large dimension with respect to the size of the gauge group. In fact, three generation of the SM fermions can be obtained from a single fermion belonging to the spinorial representation by considering some simple dynamical assumptions. Our strategy of the model building is as follows. We start from a 6D SO gauge theory with a single fermion in a spinorial representation. By compactifying the sixth spatial dimension on an orbifold $S^{1}/Z_{2}$, we obtain a model of 5D gauge-Higgs unification incorporating three generation structure.
Let us consider a 6D $SO(20)$ gauge theory with a single fermion in a spinorial representation ${\bf 1024}$, which is compactified on an orbifold $S^{1}/Z_{2}$ in a sixth dimension with a radius $R_{6}$.
@@
where $\eta_{i}$ is a sign ambiguity ±$\pm 1$. Now, we take $(\eta_{0},\eta_{1})=(+,+)$ so that the four components of ${\bf 512}_{+}$ have $+1$ eigenvalues of $i\Gamma^{6}$ and remain to be massless. In the symmetry breaking $SO(20)\to SO(11)\times SU(4)\times U(1)$ by $P_{0}$ and $P_{1}$, ${\bf 512}_{+}$ is decomposed into the following irreducible representations [26, 27].
@@
where the representations in the parentheses are those of $SO(11)\times SO(8)$ in the first line, $SO(11)\times SU(4)\times U(1)$ in the second line and the number in the subscript implies a $U(1)$ charge. The subscripts $s,c$ of ${\bf 8}$ mean the spinor and conjugated spinor representations of $SO(8)$ [26, 27].
If we assume that a subgroup $Sp(4)$ of $SU(4)$ is a confining gauge group as QCD in the SM [6], the corresponding decomposition of spinors into the irreducible representations of $Sp(4)$ are as follows [26, 27].
@@
The $Z_{2}$ parity assignments for the various components in 6D $SO(20)$ gauge field are summarized in Table 1. We can see from this table that the gauge fields of $SO(11)\times SU(4)\times U(1)$ only have $Z_{2}$ parity (+,+)(+,+)( + , + ), which means that they have massless modes. Furthermore, we have two massless mode from the sixth component of the gauge field $A_{6}$, $({\bf 11},\overline{{\bf 4}}),({\bf 1},\overline{{\bf 4}})$ since the $Z_{2}$ parity is totally opposite to those of the 5D gauge field (2). However, these fields are non-singlet in a confining subgroup $Sp(4)$ in $SU(4)$. After all, no massless scalar field comes from $A_{6}$. Note that 5D gauge field $({\bf 1},{\bf 15}_{0})$ also confines and decoupled since no $Sp(4)$ singlet is included in ${\bf 15}$ of $SU(4)$ as can be seen from the decomposition of ${\bf 15}$ of $SU(4)$ into $Sp(4)$ irreducible representations ${\bf 15}={\bf 10}+{\bf 5}$ [26, 27]. To summarize, the massless fields after compactification in the sixth dimension are the 5D $SO(11)\times U(1)$ gauge fields and three spinors in ${\bf 32}$ representation corresponding to three generations of the SM fermions.
@@
The $Z_{2}$ parity assignments for the components in 5D $SO(11)$ gauge field is summarized in Table 2. We can see from the table that the gauge fields of $SO(6)\times SO(4)$ having $Z_{2}$ parity (+,+)(+,+)( + , + ) are massless. On the other hand, $({\bf 1},{\bf 4})$ from $A_{5}$ are found to be massless and identified with the SM Higgs field. As for the fermion ${\bf 32}$, we can obtain three massless ${\bf 16}$ by following [28].
@@
where the scalar field with subscript $H$ means that it includes the SM Higgs field. Also, the symbol $\supset$ implies that it extracts zero mode terms after compactification. In the first (second) line, the gauge coupling of 6D $SO(20)$ theory is decomposed into 5D $SO(11)$(4D $SO(10)$) symmetry. Since the gauge coupling is universal in flavor space, the fermion mass hierarchy, flavor mixing and CP violation cannot be explained as it stands. Improving these points is very challenging problem and requires nontrivial extensions.
@@
There are issues to be addressed to discuss phenomenology based on this model. In family unified GHU models, Yukawa coupling constants are universal in flavor space since they are originated from the gauge coupling constant. How the fermion mass hierarchy, flavor mixing and CP violation are realized in our model is a common and important problem to be solved in models of family unification. The issues of gauge coupling unification and proton decay analysis are also important and interesting. In particular, identifying the main decay mode in our model is crucial. Although two step compactifications of 6D and the confining gauge dynamics of $Sp(4)$ were assumed in this paper, a model building based on a compactification on an orbifold $T^{2}/Z_{N}$ and that without any confining gauge group would be more interesting. These are left for our future work.

### v2

@@
In this letter, we propose a new simple model of family unification avoiding the above unsatisfactory points. This model is also based on the GGHU, but an orthogonal group is employed as a gauge group. As was discussed in [25], the reason of considering the orthogonal group is that a spinor representation in the orthogonal group has a very large dimension with respect to the size of the gauge group. In fact, three generation of the SM fermions can be obtained from a single fermion belonging to the spinorial representation by considering some simple dynamical assumptions. We emphasize here again that the SM Higgs field is also unified in the higher dimensional gauge field, which was not considered in [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Our strategy of the model building is as follows. We start from a 6D SO gauge theory with a single fermion in a spinorial representation. By compactifying the sixth spatial dimension on an orbifold $S^{1}/Z_{2}$, we obtain a model of 5D gauge-Higgs unification incorporating three generation structure.
Let us consider a 6D $SO(20)$ gauge theory with a single fermion in a spinorial representation $\bf 1024$}, which is compactified on an orbifold $S^{1}/Z_{2}$ in a sixth dimension with a radius $R_{6}$. In this model, we have no perturbative and nonperturbative gauge anomalies since 6D spinor in $\bf 1024$} is Dirac for the perturbative anomaly and the sixth homotopy group of $SO(20)$ is zero $\Pi_{6}(SO(20))=0$ for the nonperturbative anomaly.
@@
where $\eta_{i}$ is a sign ambiguity $\pm 1$. Now, we take $(\eta_{0},\eta_{1})=(+,+)$ so that the four components of $\bf 512$$_{+}$ have +1+1 eigenvalues of $i\Gamma^{6}$ and remain to be massless. In the symmetry breaking $SO(20)\to SO(11)\times SU(4)\times U(1)$ by $P_{0}$ and $P_{1}$, $\bf 512$$_{+}$ is decomposed into the following irreducible representations [26, 27].
@@
where the representations in the parentheses are those of $SO(11)\times SO(8)$ in the first line, $SO(11)\times SU(4)\times U(1)$ in the second line and the number in the subscript implies a $U(1)$ charge. The subscripts $s,c$ of $\bf 8$} mean the spinor and conjugated spinor representations of $SO(8)$ [26, 27] [^1] In general, the $2^{n-1}$ dimensional (conjugated) spinor of $SO(2n)$ can be decomposed into the sum of the (odd) even number rank totally anti-symmetric tensors on $SU(n)$, respevtively. In our case, the branching rules of 8 dimensional spinor and conjugated spinor representations under $SU(4)\times U(1)$ are $\bm{8}_{s}=[0]+[2]+[4]=\bm{1}_{2}+\bm{6}_{0}+\bm{1}_{-2},\bm{8}_{c}=[1]+[3]=\bm{4}_{-1}+\bm{\overline{4}}_{1}$, where $[n]$ represents the $n$-rank totally anti-symmetric tensor of $SU(4)$. See the page 204 of [27], for instance..
[…]
@@
The $Z_{2}$ parity assignments for the various components in 6D $SO(20)$ gauge field are summarized in Table 1. We can see from this table that the gauge fields of $SO(11)\times SU(4)\times U(1)$ only have $Z_{2}$ parity $(+,+)$, which means that they have massless modes. Furthermore, we have two massless mode from the sixth component of the gauge field $A_{6}$, $({\bf 11},\overline{{\bf 4}}),({\bf 1},\overline{{\bf 4}})$ since the $Z_{2}$ parity is totally opposite to those of the 5D gauge field (2). However, these fields are non-singlet in a confining subgroup $Sp(4)$ in $SU(4)$. After all, no massless scalar field comes from $A_{6}$. Note that 5D gauge field $({\bf 1},{\bf 15}_{0})$ also confines and decoupled since no $Sp(4)$ singlet is included in $\bf 15$} of $SU(4)$ as can be seen from the decomposition of $\bf 15$} of $SU(4)$ into $Sp(4)$ irreducible representations ${\bf 15}={\bf 10}+{\bf 5}$ [26, 27]. To summarize, the massless fields after compactification in the sixth dimension are the 5D $SO(11)\times U(1)$ gauge fields and three spinors in $\bf 32$} representation corresponding to three generations of the SM fermions.
@@
The $Z_{2}$ parity assignments for the components in 5D $SO(11)$ gauge field is summarized in Table 2. We can see from the table that the gauge fields of $SO(6)\times SO(4)$ having $Z_{2}$ parity $(+,+)$ are massless. On the other hand, $({\bf 1},{\bf 4})$ from $A_{5}$ are found to be massless and identified with the SM Higgs field. As for the fermion $\bf 32$}, we can obtain three massless $\bf 16$} by following [28].
@@
where the scalar field with subscript $H$ means that it includes the SM Higgs field [^5] Here the $SO(10)$ expression does not mean that the 4D unbroken gauge symmetry is $SO(10)$. This is just an illustration of Yukawa coupling in terms of $SO(10)$ language.. Also, the symbol $\supset$ implies that it extracts zero mode terms after compactification. In the first (second) line, the gauge coupling of 6D $SO(20)$ theory is decomposed into 5D SO(11)SO(11)(4D $SO(10)$) symmetry. Since the gauge coupling is universal in flavor space, the fermion mass hierarchy, flavor mixing and CP violation cannot be explained as it stands. Improving these points is very challenging problem and requires nontrivial extensions.
@@
There are issues to be addressed to discuss phenomenology based on this model. In family unified GHU models, Yukawa coupling constants are universal in flavor space since they are originated from the gauge coupling constant. How the fermion mass hierarchy, flavor mixing and CP violation are realized in our model is a common and important problem to be solved in models of family unification. One of the possible ways is to introduce the brane localized Yukawa interactions, which can provide the source of nontrivial flavor structure since the localized interactions does not need to respect the original symmetry. It is also useful to introduce the $Z_{2}$ odd bulk masses, which makes the mode functions with different chiralities different and provides 4D Yukawa couplings with the suppression factor by the overlap integral of the mode functions. It will be interesting to investigate the fermion flavor structure in such bulk and brane Yukawa interaction systems.
The issues of gauge coupling unification and proton decay analysis are also important and interesting. In particular, identifying the main decay mode in our model is crucial. In the model where the 6D bulk fields respecting the symmetry in the 6D bulk are introduced, the differences of the 1-loop gauge coupling running are expected to be almost logarithmic shown in [29], the unification might be realized at around $10^{14}$ GeV or so. Off course, we need to study it in detail since our model is two step breaking. In our model, the proton decay will occur via dimension six operators of $X,Y$ gauge boson and $A_{5,6}$ scalar field exchange. If the unification scale is around $10^{14}$ GeV as mentioned above, the suppression of the proton decay is not enough. However, we guess that some suppression factors come from the overlap integral of mode functions in deriving dimension six operators. If the wave functions of quarks and leptons are separately localized in extra spaces, the proton decay can be easily suppressed by the overlap integral of the wave functions [30]. Therefore, we expect that the proton decay constraints can be satisfied. In a case that the effects of the localization are not enough, it might be possible to suppress them by some symmetries (such as the discrete symmetry discussed in [31], for instance).
Although two step compactifications of 6D and the confining gauge dynamics of $Sp(4)$ were assumed in this paper, a model building based on a compactification on an orbifold $T^{2}/Z_{N}$ and that without any confining gauge group would be more interesting. These are left for our future work.
Acknowledgments
This work was supported in part by JSPS KAKENHI Grant-in-Aid for Scientific Research (C) No. 25K07304(NM) and Grant-in-Aid for JSPS Research Fellows (RN).